### v1

# Quick on the Uptake: Eliciting Implicit Intents from Human Demonstrations for Personalized Mobile-Use Agents

Abstract
As multimodal large language models advance rapidly, the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions from graphical user interface. To further enhance mobile-use agents, previous studies employ demonstration learning to improve mobile-use agents from human demonstrations. However, these methods focus solely on the explicit intention flows of humans (e.g., step sequences) while neglecting implicit intention flows (e.g., personal preferences), which makes it difficult to construct personalized mobile-use agents. In this work, to evaluate the Intention Alignment Rate between mobile-use agents and humans, we first collect MobileIAR, a dataset containing human-intent-aligned actions and ground-truth actions. This enables a comprehensive assessment of the agents’ understanding of human intent. Then we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. IFRAgent then leverages a SOP extractor combined with retrieval-augmented generation and a query rewriter to generate personalized query and SOP from a raw ambiguous query, enhancing the alignment between mobile-use agents and human intent. Experimental results demonstrate that IFRAgent outperforms baselines by an average of 6.79% (32.06% relative improvement) in human intention alignment rate and improves step completion rates by an average of 5.30% (26.34% relative improvement). The codes are available at https://github.com/MadeAgents/Quick-on-the-Uptake.
Introduction
Figure 1: Comparing IFRAgent with normal mobile-use agents and existing demonstration learning methods for mobile-use agents. IFRAgent considers both explicit intention flow and implicit intention flow, enabling it to capture more personalized information, such as taste preferences.
As multimodal large language models advance rapidly (Zhang et al. 2024b; Yin et al. 2024), the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions (e.g., clicking, scrolling) from graphical user interface (GUI) (Zhang et al. 2024a; Liu et al. 2025b). To further improve capabilities of mobile-use agents, some works (Liu et al. 2025a; Verma et al. 2024) employ demonstration learning to enable mobile-use agents to know how to complete task from human demonstrations.
However, as show in Figure 1, existing demoinstration learning method for mobile-use agents focus solely explicit intention flows of humans (e.g., operational logic, step sequences) to help mobile-use agents learn how to complete task. Moreover, user instructions in the real world are often ambiguous (Cheng et al. 2025b) and user-specific, requiring mobile-use agents to understand the implicit intention flows of humans in order to align with human intentions.
There are two challenges for mobile-use agents to align with human intentions: (i) There is a lack of datasets or benchmarks that can assess the alignment level between mobile-use agents and human intentions. (ii) Fine-tuning a separate mobile-use agent for each user to create user-specific mobile-use agents is impractical.
For challenge (i), we first collect MobileIAR, a user-specific dataset that supports both English and Chinese language users, designed to assess the Intention Alignment Rate between mobile-use agents and humans. As shown in Table 1, the dataset contains 16 apps and 945 instuction spanning 7 daily scenarios. And the dataset provides both the human-intent-aligned actions and ground-truth action lists, enabling a comprehensive assessment of the alignment level between mobile-use agents and human intentions.
For challenge (ii), we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. The IFRAgent consists of the intention flow extraction phase and the deployment phase. In the intention flow extraction phase, IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. In the deployment phase, IFRAgent, as a plug-and-play module, leverages a SOP extractor combined with retrieval-augmented generation (RAG) (Lewis et al. 2020) and a warmed-up query rewriter to rewrite the user’s ambiguous query into a user-specific personalized query and a personalized SOP based on the analysis of the human intention flow from the previous phase, thereby enabling mobile-use agents to align with human intentions.
Extensive experiments spanning diverse mobile-use agents (supporting multiple user languages and are constructed with varying methods) demonstrate that IFRAgent outperforms baseline methods by an average of 6.79% (32.06% relative improvement) in intent alignment rate and achieves a 5.30% in task completion rate (26.34% relative improvement). We also find that general-domain models (e.g., Qwen2.5-VL-7B, GPT-4o) demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS.
We further validate IFRAgent’s capability and generalizability through ablation study, cross-dataset tests, comparative studies with other methods, and scale analysis.
In summary, we make three key contributions:
(i) We contribute and open-source MobileIAR, a dataset containing both user-intent-aligned actions and ground-truth action lists. This dataset not only reflects traditional metrics such as task success rate but also measures the alignment between mobile-use agents and user intent. It establishes the first benchmark for user-specific intent alignment testing in the field of mobile-use agents.
(ii) We propose IFRAgent, a plug-and-play framework that leverages both explicit and implicit intention flows from human demonstrations to enhance mobile-use agents’ task completion capability and user-specific intent alignment.
(iii) Through extensive experiments on different mobile-use agents, we demonstrate that IFRAgent achieves an average improvement of 6.79% (32.06% relative improvement) in intent alignment rate and 5.30% in step-wise success rate (26.34% relative improvement) over baseline methods.
Dataset # Inst. # Apps # Step HL GT FS US
PixelHelp 187 4 4.2
MoTIF 276 125 4.5
UIBert 16,660 - 1
Meta-GUI 1,125 11 15
UGIF 523 12 6.3
AITW 30,378 357 6.5
AITZ 2,504 70 7.5
AndroidControl 15,283 833 4.8
AMEX 2,946 110 12.8
OS-Kairos 1000 12 5.1
MobileAgentBench 100 10 -
AppAgent 50 10 -
LlamaTouch 496 57 7.0
AndroidWorld 116 20 -
AndroidLab 138 9 8.5
LearnGUI 2353 73 13.2
IFRAgent 945 16 7.7
Table 1: Comparative analysis of datasets and environments for evaluating mobile-use agents. Key metrics include: # Inst. (instruction count), # Apps (application count), # Step (average steps per task), HL (high-level instructions), GT (ground truth trajectories), FS (few-shot learning support), and US (user-specific demonstrations). Data for this table is partially sourced from Liu et al. (2025a).
Related work
Mobile-use agents for mobile automation
Mobile-use agents (Wang et al. 2024b; Hu et al. 2025) can automatically perform mobile tasks by using the GUI to simulate human interactions (e.g., clicking, scrolling). Researchers mainly adopt two approaches to build mobile-use agents: using open-source models or using closed-source models. Some construct mobile-use (M)LLMs with open-source models  (Zhang and Zhang 2024; Hong et al. 2024; Wu et al. 2025; Xu et al. 2024b; Qin et al. 2025), often followed by post-training via reinforcement learning  (Wang et al. 2024c; Zhou et al. 2024; Luo et al. 2025; Gu et al. 2025). Others rely on the general-domain knowledge of closed-source models to develop mobile-use agents  (Jiang et al. 2025; Wang et al. 2025; Li et al. 2025). Both approaches have demonstrated strong performance in mobile automation tasks. As shown in Table 1, there are many benchmarks for evaluating mobile-use agents. These benchmarks mainly fall into two types: those that use static images for testing (Shaw et al. 2023; Li et al. 2024; Rawles et al. 2023; Zhang et al. 2024c) and those that employ dynamic environments to evaluate (Wang et al. 2024a; Zhang et al. 2025; Rawles et al. 2025; Xu et al. 2024a). However, these benchmarks only assess whether mobile-use agents complete tasks, lacking a dataset that can assess the alignment level between mobile-use agents and human intentions.
Figure 2: The workflow pipeline of IFRAgent. Above the dashed line shows the intention flow extraction phase. Below the dashed line shows the deployment phase.
Demonstration learning for mobile-use agents
Demonstration learning (Correia and Alexandre 2024) is a method that enhances the capabilities of agents by observing human demonstrations, primarily including imitation learning (Rybski et al. 2007) and inverse reinforcement learning (Ng, Russell et al. 2000). This approach has been widely applied in the robot domain (Argall et al. 2009), and in the field of mobile-use and computer-use agents, some works have utilized demonstration learning to improve agents’ ability to accomplish tasks through few-shot examples (Liu et al. 2025a; Verma et al. 2024). However, existing demonstration learning research in the mobile-use agent domain has focused solely on the explicit intention flow in human demonstrations while neglecting implicit intention flows such as user behavioral preferences. IFRAgent comprehensively analyzes both the explicit and implicit intention flows in human demonstrations, thereby establishing a more user-intent-aligned paradigm for mobile-use agent demonstration learning.
IFRAgent Framework
We propose IFRAgent, a framework that enhances intent alignment between mobile-use agents and humans. Figure 2 shows an overview. In this section, we first introduce the pipeline framework and then introduce the trainable scheme.
IFRAgent pipline
The IFRAgent pipeline can be divided into 2 phases: the intention flow extraction phase and the deployment phase. Their algorithmic representations are shown in Appendix.
Intention Flow Extraction Phase
In the intention flow extraction phase, IFRAgent collects human demonstrations and analyze these human demonstrations to extract the implicit intention flow and explicit intention flow of humans. For each user $u_{i}\in U=\{u_{1},u_{2},\dots,u_{n}\}$, we first collect human demonstrations comprising a set of queries $Q_{i}=\{q_{1},q_{2},\dots,q_{k}\}$ and initialize an empty user-level habit repository $h_{i}$. Each query $q_{j}\in Q_{i}$ is accompanied by a sequence of operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\dots,s_{p}\}$ provided by the user.
Given the tuple $(q_{j},S(u_{i},q_{j}))$, we first process it through an explicit intention flow agent $A_{e}$ to extract a SOP:
$p_{j}=A_{e}(q_{j},S(u_{i},q_{j})).$ (1)
Concurrently, the query $q_{j}$ is encoded into a latent representation $\mathbf{l}_{j}$ using an embedding model $\phi$, such that $\mathbf{l}_{j}=\phi(q_{j})$. The pair $(\mathbf{l}_{j},p_{j})$ is stored for $u_{i}$ to facilitate retrieval during deployment.
Simultaneously, the tuple is processed through an implicit intention flow agent $A_{i}$ that incrementally updates the habit repository:
$h_{i}=h_{i}+A_{i}(h_{i},q_{j},S(u_{i},q_{j})),$ (2)
where $A_{i}$ learns latent behavioral patterns from interaction sequences. This dual-processing framework iterates over all queries in $Q_{i}$ until all human demonstrations are consumed, resulting in a comprehensive habit repository $h_{i}$ and retrievable explicit SOPs $\{(\mathbf{l}_{j},p_{j})\}$ for each user.
Deployment Phase
In the deployment phase, when processing a user query $q$ from user $u_{i}$, we first encode it into vector $\mathbf{l}$ using the same embedding model $\phi$ as employed in the intention flow extraction phase (where $\mathbf{l}$ = $\phi(q)$). Then we match $\mathbf{l}$ against each explicit intention flow $(\mathbf{l}_{j},p_{j})$ of the user $u_{i}$ for RAG.
When the similarity exceeds a threshold $\tau$, we obtain the most similar query $q^{\prime}$ and its corresponding SOP $p^{\prime}$:
$(q^{\prime},p^{\prime})=\begin{cases}(q_{j},p_{j})&\text{if }\exists j=\arg\max_{k}\text{sim}(\mathbf{l},\mathbf{l}_{k})\\ &\quad\land\ \text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau,\\ \varnothing&\text{otherwise.}\end{cases}$ (3)
The query $q^{\prime}$, its SOP $p^{\prime}$, and the query $q$ are used together as the prompt for few-shot learning and then fed into the SOP Extractor $\mathcal{E}$ to obtain the SOP $p$ corresponding to $q$:
$p=\mathcal{E}(q,(q^{\prime},p^{\prime})).$ (4)
Next, the query $q$ and its corresponding SOP $p$ are combined with the user habit repository $h_{i}$ of user $u_{i}$ as input to the query writer $\mathcal{W}$ to generate a rewritten personalized query $\hat{q}$ and SOP $\hat{p}$ that align user-specific intention:
$(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i}).$ (5)
Finally, the rewritten query $\hat{q}$, the rewritten SOP $\hat{p}$, and the current screenshot $s$ are provided as input to the mobile-use agent $\mathcal{F}$ to obtain the action $a$:
$a=\mathcal{F}(\hat{q},\hat{p},s).$ (6)
Trainable Scheme
The SOP Extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are two key components of the IFRAgent during the deployment phase. Both $\mathcal{E}$ and $\mathcal{W}$ are models deployable on the edge side, thus lacking general knowledge about mobile operations. Since $\mathcal{E}$ has already compensated for the lack of general knowledge through few-shot learning with RAG, so only $\mathcal{W}$ require supervised fine-tuning to unleash its potential in query rewriting.
To warm up the trainable scheme query writer $\mathcal{W}$, we first employ a crowdsourcing approach to directly collect data from a group of English and Chinese users, who fill out the habit repository $H=\{h_{1},h_{2},\dots,h_{n}\}$. Then, using a combination of manual construction and LLM-generated expansions, we create a set of ambiguous instructions $q$. These users are asked to provide their own SOP $p$ for each ambiguous instruction, as well as their personalized rewritten queries $\hat{q}$ and personalized rewritten SOPs $\hat{p}$.
Next, we train the query writer $\mathcal{W}$ to predict the personalized rewritten query $\hat{q}$ and the personalized rewritten SOP $\hat{p}$, given the user $u_{i}$’s habit repository $h_{i}$, the ambiguous instruction $q$, and the corresponding SOP $p$. We use this dataset to train the query writer $\mathcal{W}$ for one epoch to achieve the warm-up effect.
The training objective can be formulated as:
[…]
After warm-up training, $\mathcal{W}$ acquires the capability to rewrite queries and SOPs in a user-specific manner.
Experiments
English users Chinese users
[…]
Open-source based mobile-use agents
OS-Atlas-7B-Pro 42.30 75.39 36.65 42.22 76.92 35.67
                    +IFRAagent 48.466.16↑ 80.985.59↑ 43.466.81↑ 51.979.75↑ 81.644.72↑ 47.8212.15↑
UI-TARS-7B-SFT 43.11 69.26 37.28 44.86 75.00 36.86
                    +IFRAagent 44.481.37↑ 72.573.31↑ 40.693.41↑ 51.046.18↑ 75.860.86↑ 48.7011.84↑
UI-TARS-7B-DPO 41.12 71.11 34.96 45.07 70.74 37.19
                    +IFRAagent 41.580.46↑ 70.910.20↓ 37.732.77↑ 46.451.38↑ 73.732.99↑ 43.466.27↑
UI-TARS-1.5-7B 40.19 72.06 34.02 46.22 76.42 39.18
                    +IFRAagent 42.782.59↑ 72.720.66↑ 39.265.24↑ 49.803.58↑ 77.010.59↑ 46.757.57↑
Qwen2.5-VL-7B 12.29 16.13 11.80 19.29 23.52 18.13
                    +IFRAagent 30.5718.28↑ 40.6724.54↑ 27.7015.90↑ 38.2718.98↑ 47.2723.75↑ 36.1318.00↑
Close-source based mobile-use agents
GPT-4o+OCR model 35.63 74.75 32.05 37.13 77.42 31.18
                    +IFRAagent 40.554.92↑ 74.500.25↓ 36.023.97↑ 44.197.06↑ 78.060.64↑ 41.4010.22↑
GLM-4v+OCR model 2.54 57.94 1.76 3.11 72.97 2.47
                    +IFRAagent 3.210.67↑ 73.1415.20↑ 2.470.71↑ 4.050.94↑ 73.710.74↑ 3.030.56↑
Qwen-VL-max+OCR model 19.86 79.91 16.69 22.00 81.55 19.04
                    +IFRAagent 20.370.51↑ 81.821.91↑ 17.560.87↑ 24.042.04↑ 84.422.87↑ 21.342.30↑
Table 2: Performance comparison of mobile-use agents with IFRAagent enhancements, categorized by open-source and close-source models. IFRAgent demonstrates improvements across nearly all metrics.
Experiments Setup
Baseline
Mobile-use agents can be categorized by their construction methods into open-source and closed-source agents. We selected five state-of-the-art open-source based mobile-use agents and three close-source based mobile-use agents as baselines to experimentally validate the extent to which IFRAgent enhances mobile-use agents. The open-source based mobile-use agents include OS-Atlas-7B-Pro (Wu et al. 2025), UI-TARS-7B-SFT, UI-TARS-7B-DPO (Qin et al. 2025), UI-TARS-1.5-7B (Seed 2025), and Qwen2.5-VL-7B (Bai et al. 2025), while the close-source based mobile-use agents include GPT-4o (OpenAI 2023), GLM-4v (GLM et al. 2024), and Qwen-VL-max (Bai et al. 2023). Since GPT-4o, GLM-4v, and Qwen-VL-max inherently lack the capability to predict coordinates, we incorporated an OCR model composed of ResNet18 (He et al. 2016) and ConvNeXt-Tiny (Liu et al. 2022) to assist them in localization. Our main experiments were conducted on our collected dataset, MobileIAR. MobileIAR is the first benchmark for user-specific intent alignment testing in the field of mobile-use agents. In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP extractor and query rewriter are based on Qwen3-4B.
MobileIAR Dataset collection
We collect trajectories for English users and Chinese users in 7 daily life scenarios through a crowdsourcing approach, resulting in a total of 945 instructions and 7,310 screenshots. For each user in each daily life scenario, there are 5 trajectories consisting only of queries and screenshots as the support dataset, and 10 trajectories containing queries, screenshots, human-intent-aligned actions, and ground-truth action lists as the test dataset. Both the human-intent-aligned actions and ground-truth action lists are manually annotated by users. The human-intent-aligned action is the single action that best aligns with human intent for the current query and screenshot. The ground-truth action list includes the human-intent-aligned action while potentially containing other correct but less optimal actions—for example, selecting a dish that does not match the user’s preferred flavor when ordering food. The support dataset serves as human demonstrations for the IFRAgent to extract intention flows, while the test dataset is used for metric calculations.
Metric
We consider two types of metrics: one measures the task completion capability of mobile-use agents, and the other measures the alignment level between mobile-use agents and human intentions. Following existing work on mobile-use agents (Zhang and Zhang 2024; Ma, Zhang, and Zhao 2024; Wu et al. 2025; Qin et al. 2025; Cheng et al. 2025a), we report the step-wise success rate (SR) and action type accuracy (Type) to assess task completion. To measure the alignment level between mobile-use agents and human intentions, we report the intention alignment rate (IAR). In the MobileIAR dataset, we provide both human-intent-aligned actions and ground-truth action lists. The human-intent-aligned action is the single most aligned action with the user’s intent at each step, while ground-truth action lists are a set of possible actions that could help fulfill the user’s query in the current frame. For metric calculations, SR and Type consider the mobile-use agent’s action correct if it matches any of the ground-truth actions at the current step. In contrast, IAR requires the agent’s action to exactly match the human-intent-aligned action to be counted as correct.
Main Results
As shown in Table 2, we conducted extensive experiments on mobile-use agents constructed using 8 different methods. Based on these results, we have the following key findings:
(i) IFRAgent can improve almost all mobile-use agents, with absolute improvements of SR by 5.30% and IAR by 6.79%, and relative improvements of SR by 26.34% and IAR by 32.06%. This indicates that IFRAgent can enhance both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. Moreover, these improvements are effective for both English users and Chinese users.
(ii) Among general-domain models, Qwen2.5-VL-7B and GPT-4o demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS. This is because models with broader general knowledge tend to have a more accurate and comprehensive understanding of human intentions. Specialized mobile-use agent base models, due to the extensive GUI operation knowledge learned during the post-training process, tend to forget some general world knowledge that is helpful for human intention recognition.
(iii) Building personalized mobile-use agents requires the model to possess basic instruction-following capabilities. We can observe that the Type metric of mobile-use agents built using GLM-4v and Qwen-VL-max is significantly lower than the SR metric, indicating that in most cases, they can determine that the current action type required is CLICK but fail to utilize the localization information provided by the OCR model. Therefore, even with the addition of the IFRAgent module, the improvement for mobile-use agents built with GLM-4v and Qwen-VL-max is not significant, as IFRAgent does not provide localization information for the mobile-use agent.
Overall, IFRAgent effectively enhances mobile-use agents by improving both task completion and human intention understanding, particularly for general-domain models.
Further Analysis
Ablation Study
Figure 3: Experimental results of ablation study.
The SOP extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are the two most critical components in IFRAgent. We conducted an ablation study on Qwen2.5-VL-7B, the model with the most significant improvement from IFRAgent. The experimental results are shown in Figure 3. It can be observed that the SOP extractor $\mathcal{E}$ alone cannot fully unleash the potential of human demonstrations; after personalized rewriting, all metrics still show improvement. On the other hand, the query writer $\mathcal{W}$ alone, which only rewrites abstract personalized queries without SOP teaching the mobile-use agent how to act, can only bring slight improvements over the baseline. Therefore, the design of IFRAgent maximizes the potential mined from human demonstrations and is reasonable.
Generalizability Analysis on Other Datasets
[…]
OS-Atlas-7B-Pro 58.85 79.52 53.78
                    +IFRAagent 69.74 85.31 68.42
UI-TARS-1.5-7B 61.49 75.86 53.24
                    +IFRAagent 60.75 77.04 58.80
Qwen2.5-VL-7B 24.27 27.77 23.64
                    +IFRAagent 51.64 59.62 50.31
GPT-4o+OCR model 40.65 68.92 37.67
                    +IFRAagent 56.33 77.36 53.75
Table 3: Generalizability experiment on our modified user-specific OS-Kairos dataset.
To validate the generalizability of the IFRAgent method, we adapt a portion of the OS-Kairos dataset (Cheng et al. 2025a). We simulate the extraction of trajectories across shop, video and search scenarios in the training set of OS-Kairos as human demonstrations. Then, in the corresponding scenarios of the OS-Kairos test set, we sample user trajectories and manually annotate the user-intent-aligned actions and ground-truth action lists.
The experimental results are presented in Table 3. IFRAgent continues to demonstrate improvements in both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios on our modified user-specific OS-Kairos dataset, proving the generalizability of IFRAgent.
At the same time, consistent with the conclusions in the main results, general-domain models such as Qwen2.5-VL-7B and GPT-4o still show better performance than specialized mobile-use agent base models like UI-TARS.
Figure 4: Experimental results of IFRAgent on model scale for Qwen2.5-VL-3B, Qwen2.5-VL-7B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. The improvement is more significant for medium-scale mobile-use agents.
Comparison With Other Methods
[…]
OS-Atlas-7B-Pro 42.26 76.24 36.11
           +SOP demonstration 50.08 79.65 43.72
           +IFRAagent 50.40 81.35 45.87
UI-TARS-1.5-7B 43.53 74.48 36.88
           +SOP demonstration 41.32 70.86 37.17
           +IFRAagent 46.67 75.10 43.41
Qwen2.5-VL-7B 16.17 20.22 15.31
           +SOP demonstration 22.64 28.68 21.11
           +IFRAagent 34.83 44.32 32.37
GPT-4o+OCR model 36.46 76.23 31.57
           +SOP demonstration 43.37 76.13 39.29
           +IFRAagent 42.57 76.47 39.00
Table 4: Comparison with SOP demonstration. IFRAgent can better abstract the user’s intention flow, thereby significantly enhancing the agent.
Implementation Details
To validate the effectiveness of IFRAgent in recognizing intention flow from human demonstrations, we compared it with other methods that extract information from human demonstrations. Since Learnact (Liu et al. 2025a) is not a user-specific solution, it cannot be directly tested on MobileIAR. Instead, we emulated Learnact’s approach by using a demoparser to extract the SOP from a set of user-specific human demonstrations. For each query, we matched it with the corresponding 1-shot SOP demonstration to enhance the prompt. We then selected four representative baselines on MobileIAR to compare this SOP demonstration method with IFRAgent. Finally, we rewrite all the queries into ambiguous instructions to facilitate testing for IAR.
Result Analysis
As shown in Table 4, the experimental results demonstrate that all demonstration learning methods can improve various metrics across different baselines. However, for almost all baselines, IFRAgent achieves more significant improvements compared to SOP demonstration.
For the mobile-use agent built with GPT-4o and the OCR model, the performance of IFRAgent and SOP demonstration is relatively close. This is because GPT-4o, as a foundation model with extensive world knowledge, can leverage its own capabilities to abstract the user’s implicit intent flow from the SOP demonstration to assist in task completion.
For most 7B-parameter-scale agents, whether general-domain models or specialized mobile-use agent base models, IFRAgent shows significantly better performance than SOP demonstration. This is because 7B-scale agents struggle to accurately abstract implicit intent flow from SOP demonstrations. In contrast, IFRAgent already incorporates the user’s explicit and implicit intent flows into the personalized rewritten query and personalized rewritten SOP, eliminating the need for the agent to perform further abstraction.
Scale Analysis
Model scale Analysis
To analyze the impact of IFRAgent on mobile-use agents with different parameter scales and simultaneously verify its generalizability across varying model sizes, we conducted experiments on Qwen2.5-VL-3B, Qwen2.5-VL-7B, and Qwen2.5-VL-72B, respectively. As shown in Figures 4, IFRAgent improves both task completion capability and alignment with human intentions for mobile-use agents of different model scales. The most significant improvements are observed in the 7B and 32B model scales. This is because the 3B-parameter mobile-use agent lacks sufficient instruction-following ability—even when provided with personalized queries and SOPs, the mobile-use agent still fails to comply. On the other hand, the 72B-parameter mobile-use agent already possesses strong general mobile operation capabilities, so the SR improvement is less pronounced. However, there remains a noticeable enhancement in the IAR for the 72B-parameter mobile-use agent. IFRAgent stimulates both the general task completion capability and the ability to align with human intentions in mobile-use agents with moderate parameter scales.
Figure 5: Experimental results of IFRAgent-enhanced model under varying numbers of demonstrations.
Demonstration Count Analysis
To analyze the impact of the number of demonstration queries and SOPs on the Extractor $\mathcal{E}$, we conducted a demonstration count analysis on Qwen2.5-VL-7B. We input varying numbers of $(q^{\prime},p^{\prime})$ pairs into $\mathcal{E}$ for SOP extraction. The experimental results are shown in Figure 5. The results show that while 1-shot brings a significant improvement over 0-shot, increasing to 2-shot and 3-shot does not lead to further substantial gains across metrics and may even introduce irrelevant information that interferes with $\mathcal{E}$, causing performance degradation. On one hand, this indicates that the intent flow extracted by IFRAgent from human demonstrations can indeed help mobile-use agents align with human intent. On the other hand, increasing the number of demonstrations entails higher computational overhead. Therefore, using 1-shot for SOP extraction in IFRAgent is a reasonable choice.
Conclusion
This study focuses on two major challenges in building personalized mobile-use agents: the lack of alignment evaluation benchmarks and the impracticality of per-user fine-tuning. For these challenges, this study collects and open-source the MobileIAR dataset, a user-specific dataset designed to assess the intention alignment rate between mobile-use agents and humans. This study then proposes IFRAgent, a plug-and-play framework based on intention flow recognition from human demonstrations, which enhances the alignment between mobile-use agents and human intentions by leveraging human demonstrations to rewrite vague user instructions into personalized queries and SOPs. The extensive experiment results demonstrate that IFRAgent improves both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. This study also provides various analysis to offer valuable insights for building personalized mobile-use agents.
Appendix
Appendix A Ethics Statement
The MobileIAR dataset we collected has undergone obscuration processing for all involved personal sensitive information, ensuring no security impact on anyone’s privacy. All models used in this work are sourced from the official repositories associated with the original papers, and we strictly follow their respective usage protocols. Furthermore, we have explicitly cited them in the main text.
Appendix B Baselines details
Qwen2.5-VL
The Qwen2.5-VL series includes models with parameters of 3B, 7B, 32B, and 72B, significantly enhancing capabilities in multitask scenarios such as document parsing, object detection, and mathematical reasoning. The scale of its pre-training data has been greatly expanded to 4.1 trillion tokens, covering high-quality content such as text-image interactions, OCR, and video localization. Additionally, the Qwen2.5-VL series has undergone further adaptation for agent scenarios.
OS-Atlas
OS-Atlas is a specialized model in the field of GUI agents, pre-trained and fine-tuned in the domain of GUI agents. OS-Atlas provides two model sizes: 4B and 7B. The 4B version is fine-tuned from InternVL2-4B, while the 7B version is fine-tuned from Qwen2-VL-7B-Instruct.
UI-TARS
UI-TARS is a specialized model in the field of GUI agents, developed from the Qwen2-VL series through enhanced perception, unified action modeling, system-2 reasoning, and iterative training with reflective online traces. It includes parameter scales of 2B, 7B, and 72B, with both SFT and DPO versions released for each parameter scale.
UI-TARS-1.5
UI-TARS-1.5 utilizes the architecture and technology of UI-TARS, further integrating advanced reasoning enabled by reinforcement learning. This allows the model to reason through its thoughts before taking action, significantly enhancing its performance and adaptability, particularly in inference-time scaling.
GPT-4o
GPT-4o is an advanced model in the generative pre-trained transformer series, characterized by improved language understanding and generation. With enhanced fine-tuning capabilities and a larger training dataset, it serves as a strong baseline for evaluating performance in various natural language processing tasks.
GLM-4v
GLM-4V is the multimodal version of GLM-4, achieving a deep integration of visual and language features without compromising performance on any NLP tasks. It supports various image and video understanding tasks, including visual question answering, image captioning, visual localization, and complex object detection.
Appendix C Experiments details
Hardware information
The experiment was conducted on a server running Ubuntu 22.04.5 LTS. The hardware configuration includes an Intel(R) Xeon(R) Platinum 8358 CPU with a total of 128 cores (32 cores per socket across two sockets), operating at a maximum frequency of 3.4 GHz and a minimum frequency of 800 MHz, along with a total memory of 1.0 TiB. The server is equipped with eight NVIDIA A800-SXM4-80GB GPUs, utilizing driver version 550.54.14 and CUDA version 12.4. And each GPU has a total memory of 81,920 MiB.
Other experiments details
In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP Extractor and query rewriter are based on Qwen3-4B. The Qwen3-4B used as the query rewriter underwent a warm-up, with a training learning rate of 1e-6 and a training epoch of one epoch.
Appendix D Algorithm
For the IFRAgent framework, we present the algorithm for the intention flow extraction phase in Algorithm 1, and the algorithm for the deployment phase in Algorithm 2.
Algorithm 1 Intention Flow Extraction Phase Algorithm
1:Input: User set $U=\{u_{1},u_{2},\ldots,u_{n}\}$ with queries $Q_{i}$
2:Output: User explicit SOPs and implicit habit repository $h_{i}$ for each user $u_{i}\in U$.
3:for each user $u_{i}\in U$ do
4:  Initialize user-level habit repository $h_{i}$.
5:  Collect queries $Q_{i}=\{q_{1},q_{2},\ldots,q_{k}\}$ from $u_{i}$.
6:  for each query $q_{j}\in Q_{i}$ do
7:   Retrieve operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\ldots,s_{p}\}$.
8:   $p_{j}=A_{e}(q_{j},S(u_{i},q_{j}))$ // Extract SOP using explicit intention flow agent $A_{e}$
9:   $\mathbf{l}_{j}=\phi(q_{j})$ // Encode query into latent representation
10:   Store pair $(\mathbf{l}_{j},p_{j})$ in user-level repository.
11:   $h_{i}=h_{i}+A_{i}(h_{i},q_{j},S(u_{i},q_{j}))$ // Update habit repository with implicit intention flow
12:  end for
13:end for
14:return  User habit repository $h_{i}$ and explicit SOPs
Algorithm 2 Deployment Phase Algorithm
1:Input: User query $q$, user habit repository $h_{i}$, explicit SOPs $(\mathbf{l}_{k},p_{k})$
2:Output: Action $a$ for user query $q$
3: Encode the query into latent representation: $\mathbf{l}=\phi(q)$
4: Initialize $(q^{\prime},p^{\prime})\leftarrow\varnothing$.
5:for each $(\mathbf{l}_{k},p_{k})$ in stored SOPs of user $u_{i}$ do
6:  if $\text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau$ then
7:   if $\text{sim}(\mathbf{l},\mathbf{l}_{k})>\text{sim}(\mathbf{l},\mathbf{l}^{\prime})$ then
8:    $(q^{\prime},p^{\prime})\leftarrow(q_{k},p_{k})$ // Update most similar query
9:   end if
10:  end if
11:end for
12:if $(q^{\prime},p^{\prime})\neq\varnothing$ then
13:  Extract SOP for query $q$: $p=\mathcal{E}(q,(q^{\prime},p^{\prime}))$
14:  $(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i})$ // Generate personalized query and SOP
15:  $a=\mathcal{F}(\hat{q},\hat{p},s)$ // Obtain action from GUI agent
16:end if
17:return  Action $a$ for user query $q$
Appendix E Prompt
In this section, we present the prompts used in our study. During the intention flow extraction phase, the prompt for the implicit intention flow agent is shown in Figure 6, while the prompt for the explicit intention flow agent is shown in Figure 7. In the deployment phase, the prompt for the SOP extractor in the IFRAgent framework is illustrated in Figure 8, and the prompt for the query rewriter is shown in Figure 10. Additionally, the prompts involving different mobile-use agents in the experiments are displayed in Figure 12, Figure 13, Figure 9, Figure 11, Figure 14, and Figure 15.
Appendix F Dataset Construction Details of MobileIAR
MobileIAR collected approximately [number] images from 945 user execution trajectories across 16 apps covering 7 high-frequency daily scenarios. The dataset includes human demonstrations from 9 users (4 English users and 5 Chinese users). Each user contributed 15 trajectories per high-frequency scenario, with 5 trajectories serving as the support dataset for human demonstrations and 10 trajectories allocated as the test dataset for evaluation. The action space for the IFRAgent dataset is shown in Table 5.
Action Type Action Description
CLICK Click at the specified position.
TYPE Enter specified text at the designated location.
SCROLL Scroll in the specified direction.
PRESS_BACK Press a back button to navigate to the previous screen.
PRESS_HOME Press a home button to navigate to the home page.
WAIT Wait for the screen to load.
LONG_PRESS Long press at the specified position.
COMPELTE Indicate the task is finished.
Table 5: Action space for MobileIAR dataset.
MobileIAR is the first dataset in the field of mobile-use agents to provide user-specific data. For the test dataset trajectories, we not only include the conventional ground truth actions typically found in mobile-use or computer-use agents, but also introduce a dual annotation framework that accounts for individual user preferences. Specifically, we provide both the action most aligned with human intent (to evaluate agent-user alignment) and a set of task-beneficial actions (to assess general task completion capability). This approach enables comprehensive evaluation of mobile-use agents across both user-centric and task-oriented dimensions.
Appendix G Metric Definitions Details
In our experiments, we primarily report three metrics: step-wise success rate (SR), action type accuracy (Type), and intention alignment rate (IAR). The specific calculation details are as follows:
For each screenshot $s_{i}$ in the test set, we annotate the most intention-aligned action $a_{i}$ and a set of equally correct but not the most intention-aligned actions $A_{i}=\{a_{i1},a_{i2},...,a_{in}\}$, where $a_{i}\in A_{i}$. When the mobile user agent generates an action $a$ for the current screenshot $s_{i}$ and instruction query $q_{i}$: If $a$ matches any element in $A_{i}$, SR is incremented by 1. If the action type of $a$ matches the action type of any element in $A_{i}$, Type is incremented by 1. If $a$ matches $a_{i}$, IAR is incremented by 1 (finally, these counts are divided by the total number of samples—please help me formalize this part).
Here, CLICK and LONG_PRESS require a relative error of less than 14%, and TYPE requires a text similarity of over 80%. SCROLL, PRESS_BACK, PRESS_HOME, and WAIT must be completely accurate.
Formal definition suggestion for the normalization part: Let $N$ be the total number of test samples. Then:
$\text{SR}=\frac{\sum_{i=1}^{N}\mathbb{I}(a\in A_{i})}{N},$ (8)
$\text{Type}=\frac{\sum_{i=1}^{N}\mathbb{I}(\text{type}(a)\in\{\text{type}(a^{\prime})|a^{\prime}\in A_{i}\})}{N},$ (9)
$\text{IAR}=\frac{\sum_{i=1}^{N}\mathbb{I}(a=a_{i})}{N},$ (10)
where $\mathbb{I}(\cdot)$ is the indicator function.
Thus, IAR is always less than or equal to SR, which in turn is always less than or equal to Type.
$\text{IAR}\leq\text{SR}\leq\text{Type}$ (11)
The closer IAR and SR are, the more fully the potential of mobile user agents to align with human intentions is realized. The closer SR and Type are, the more accurate the localization of mobile user agents is.
Figure 6: Implicit intention flow agent prompt.
Figure 7: Explicit intention flow agent prompt.
Figure 8: SOP extractor prompt.
Figure 9: UI-TARS and TARS-1.5 test prompt.
Figure 10: Query rewriter prompt.
Figure 11: UI-TARS and TARS-1.5 with IFRAgent test prompt.
Figure 12: OS-Atlas test prompt.
Figure 13: OS-Atlas with IFRAgent test prompt.
Figure 14: Qwen2.5-VL test prompt.
Figure 15: Qwen2.5-VL with IFRAgent test prompt.