### v2

# Quick on the Uptake: Eliciting Implicit Intents from Human Demonstrations for Personalized Mobile-Use Agents

Abstract.
As multimodal large language models advance rapidly, the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions from graphical user interfaces. To further enhance mobile-use agents, previous studies employ demonstration learning to improve mobile-use agents from human demonstrations. However, these methods focus solely on the explicit intention flows of humans (e.g., step sequences) while neglecting implicit intention flows (e.g., personal preferences), which makes it difficult to construct personalized mobile-use agents. In this work, to evaluate the Intention Alignment Rate between mobile-use agents and humans, we first collect MobileIAR, a dataset containing human-intent-aligned actions and ground-truth actions. This enables a comprehensive assessment of the agents’ understanding of human intent. Then we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. IFRAgent then leverages a SOP extractor combined with retrieval-augmented generation and a query rewriter trained through knowledge distillation to generate personalized queries and SOPs from a raw ambiguous query, enhancing the alignment between mobile-use agents and human intent. Experimental results demonstrate that IFRAgent outperforms baselines by an average of 28.80% relative improvement in human intention alignment rate and improves step completion rates by an average of 24.20% relative improvement.
GUI agent, MLLM agent, Personalized agent, Mobile-use agent
submissionid: 417copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
1. Introduction
Figure 1. Comparing IFRAgent with normal mobile-use agents and existing demonstration learning methods for mobile-use agents. IFRAgent considers both explicit intention flow and implicit intention flow, enabling it to capture more personalized information, such as taste preferences.
As multimodal large language models advance rapidly (Zhang et al., 2024b; Yin et al., 2024; Chen et al., 2026), the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions (e.g., clicking, scrolling) from graphical user interface (GUI) (Zhang et al., 2024a; Liu et al., 2025b; Tang et al., 2025b). To further improve the capabilities of mobile-use agents, some works (Liu et al., 2025a; Verma et al., 2024) employ demonstration learning to enable mobile-use agents to know how to complete tasks from human demonstrations.
However, as shown in Figure 1, existing demonstration learning methods for mobile-use agents focus solely on explicit human intention flows (e.g., operational logic, step sequences) to help mobile-use agents learn how to complete tasks. Moreover, user instructions in the real world are often ambiguous (Cheng et al., 2025b) and user-specific, requiring mobile-use agents to understand the implicit intention flows of humans in order to align with human intentions.
There are two challenges for mobile-use agents to align with human intentions: (i) There is a lack of datasets or benchmarks that can assess the alignment level between mobile-use agents and human intentions. (ii) Fine-tuning a separate mobile-use agent for each user to create user-specific mobile-use agents is impractical.
For challenge (i), we first collect MobileIAR. This user-specific dataset supports both English and Chinese language users, designed to assess the Intention Alignment Rate between mobile-use agents and humans. The dataset contains 16 apps and 945 instructions spanning seven daily scenarios. Moreover, the dataset provides both the human-intent-aligned actions and ground-truth action lists, enabling a comprehensive assessment of the alignment level between mobile-use agents and human intentions.
For challenge (ii), we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. The IFRAgent consists of the intention flow extraction phase and the deployment phase. In the intention flow extraction phase, IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. In the deployment phase, IFRAgent, as a plug-and-play module, leverages a SOP extractor combined with retrieval-augmented generation (RAG) (Lewis et al., 2020) and a knowledge distillation warmed-up query rewriter to rewrite the user’s ambiguous query into a user-specific personalized query and a personalized SOP based on the analysis of the human intention flow from the previous phase. This enables mobile-use agents to better align with human intentions, while also improving their general task completion capability to a certain extent.
Extensive experiments spanning diverse mobile-use agents (supporting multiple user languages and constructed with varying methods) demonstrate that IFRAgent outperforms baseline methods by an average of 6.79% (28.80% relative improvement) in intent alignment rate and achieves a 5.30% in task completion rate (24.20% relative improvement). We also find that general-domain models (e.g., Qwen2.5-VL, GPT-4o) demonstrate more significant improvements compared to specialized mobile-use agent base models.
We further validate IFRAgent’s capability and generalizability through an ablation study, cross-dataset tests, comparative studies with other methods, and scale analysis.
In summary, we make three key contributions:
(i) We contribute and open-source MobileIAR, a dataset containing both user-intent-aligned actions and ground-truth action lists. This dataset not only reflects traditional metrics such as task success rate but also measures the alignment between mobile-use agents and user intent. It establishes the first benchmark for user-specific intent alignment testing in the field of mobile-use agents.
(ii) We propose IFRAgent, a plug-and-play framework that leverages both explicit and implicit intention flows from human demonstrations to enhance mobile-use agents’ task completion capability and user-specific intent alignment.
(iii) Through extensive experiments on different mobile-use agents, we demonstrate that IFRAgent achieves an average improvement of 6.56% (32.06% relative improvement) in intent alignment rate and 5.50% in step-wise success rate (26.34% relative improvement) over baseline methods.
Table 1. Comparative analysis of datasets and environments for evaluating mobile-use agents. Key metrics include: # Inst. (instruction count), # Apps (application count), # Step (average steps per task), HL (high-level instructions), GT (ground truth trajectories), FS (few-shot learning support), and US (user-specific demonstrations). Data for this table is partially sourced from Liu et al. (2025a).
Dataset # Inst. # Apps # Step HL GT FS US
PixelHelp 187 4 4.2
MoTIF 276 125 4.5
UIBert 16,660 - 1
Meta-GUI 1,125 11 15
UGIF 523 12 6.3
AITW 30,378 357 6.5
AITZ 2,504 70 7.5
AndroidControl 15,283 833 4.8
AMEX 2,946 110 12.8
OS-Kairos 1000 12 5.1
MobileAgentBench 100 10 -
AppAgent 50 10 -
LlamaTouch 496 57 7.0
AndroidWorld 116 20 -
AndroidLab 138 9 8.5
LearnGUI 2353 73 13.2
MobileIAR 945 16 7.7
Figure 2. Examples of human-intent-aligned actions and SR ground truth lists. Even for the same query, the human-intent-aligned actions of user 1, user 2, and user 3 are different. However, these actions, along with some other actions that are also beneficial for completing the query, all belong to the SR ground truth list.
Figure 3. Construction of the MobileIAR dataset. In Round 1, annotators independently label human intent-aligned actions. In Round 2, groups of three participants are formed to jointly label the SR (step-wise success rate) ground truth lists based on the trajectories obtained from Round 1.
2. Related work
2.1. Mobile-use agents for mobile automation
Mobile-use agents (Wang et al., 2024b; Hu et al., 2025) can automatically perform mobile tasks by using the GUI to simulate human interactions (e.g., clicking, scrolling). Researchers mainly adopt two approaches to build mobile-use agents: using open-source models or using closed-source models. Some construct mobile-use (M)LLMs with open-source models  (Zhang and Zhang, 2024; Hong et al., 2024; Wu et al., 2025; Xu et al., 2024b; Qin et al., 2025), often followed by post-training via reinforcement learning  (Wang et al., 2024c; Zhou et al., 2024; Lu et al., 2026; Tang et al., 2025a). Others rely on the general-domain knowledge of closed-source models to develop mobile-use agents  (Jiang et al., 2025; Wang et al., 2025b; Li et al., 2025). Both approaches have demonstrated strong performance in mobile automation tasks. As shown in Table 1, there are many benchmarks for evaluating mobile-use agents. These benchmarks mainly fall into two types: those that use static images for testing (Shaw et al., 2023; Li et al., 2024; Rawles et al., 2023; Zhang et al., 2024c) and those that employ dynamic environments to evaluate (Wang et al., 2024a; Zhang et al., 2025; Rawles et al., 2025; Xu et al., 2024a). However, these benchmarks only assess whether mobile-use agents complete tasks, lacking a dataset that can assess the alignment level between mobile-use agents and human intentions.
2.2. Demonstration learning for mobile-use
Demonstration learning (Correia and Alexandre, 2024) is a method that enhances the capabilities of agents by observing human demonstrations, primarily including imitation learning (Rybski et al., 2007) and inverse reinforcement learning (Ng et al., 2000). This approach has been widely applied in the robot domain (Argall et al., 2009), and in the field of mobile-use and computer-use agents, some works have utilized demonstration learning to improve agents’ ability to accomplish tasks through few-shot examples (Liu et al., 2025a; Verma et al., 2024). Another popular approach to building mobile-use agents via demonstration learning is to leverage videos (Zhang et al., 2026; Jung et al., 2026; Wang et al., 2025a). By efficiently processing videos, valuable data can be extracted to serve as references for mobile-use agents or to synthesize training data. However, existing demonstration learning research in the mobile-use agent domain has focused solely on the explicit intention flow in human demonstrations while neglecting implicit intention flows such as user behavioral preferences. IFRAgent comprehensively analyzes both the explicit and implicit intention flows in human demonstrations, thereby establishing a more user-intent-aligned paradigm for mobile-use agent demonstration learning.
3. MobileIAR dataset collection
In this section, we first explain why the emergence of MobileIAR is necessary, and then we introduce the specific process of constructing MobileIAR.
3.1. Motivation of MobileIAR dataset
In real-world applications, the same query given by different users can have different implicit meanings. Ideal mobile-use agents need to understand the implicit intent behind the user’s query, enabling personalized operation of the user’s smartphone. Such personalized mobile-use agents will significantly enhance the user experience and drive development in the agent-human interaction paradigm.
However, in order to construct personalized mobile-use agents, the first issue to address is how to assess the level of personalization and alignment between agents and human intentions. Currently, there is a lack of datasets or benchmarks that can evaluate the alignment between mobile-use agents and human intentions. Therefore, we propose the MobileIAR dataset and provide a measurement method for the IAR (Intention Alignment Rate) metric within MobileIAR, laying the foundation for personalized mobile-use agents. As shown in Figure 2, MobileIAR provides both human-intent-aligned actions to measure the degree of intent alignment, and ground truth lists to measure task completion capability.
3.2. Construction of MobileIAR dataset
To construct MobileIAR dataset, as shown in Figure 3, we first collect trajectories for both English-speaking and Chinese-speaking users across seven daily-life scenarios through a crowdsourcing approach, resulting in a total of 945 instructions and 7,310 screenshots. The annotators cover a diverse group of participants, including men and women from different age groups.
In the annotation process, we conduct two rounds of annotation. The first round of annotation is done independently by the annotators, where they collect trajectories on their own real mobile devices. We encourage the annotators to choose completion methods that highlight their personalized preferences for the same instruction. The results of this round are used as the human-intent-aligned action. The human-intent-aligned action measures the ability of the mobile-use agent to align with the user’s intent.
The second round of annotation builds upon the first round, aiming to allow MobileIAR to assess not only the mobile-use agent’s intent alignment capability but also its general competence. Since the human-intent-aligned action is not always the only correct action in some cases, and often the agent’s execution path may have multiple routes, we assign each trajectory from the first round randomly to three annotators to label possible correct alternative actions. The intersection of these labeled actions is then taken as the ground-truth action list. The ground-truth action list measures the mobile-use agent’s general task completion ability.
4. IFRAgent Framework
Figure 4. The workflow pipeline of IFRAgent. The phase above the dashed line represents the intention flow extraction phase, while the phase below the dashed line represents the deployment phase. In the intention flow extraction phase, IFRAgent extracts SOPs and user profiles based on users’ trajectories. In the deployment phase, IFRAgent uses this information to rewrite personalized queries and generate personalized SOPs.
We propose IFRAgent, a framework that enhances intent alignment between mobile-use agents and humans. Figure 4 shows an overview. In this section, we first introduce the pipeline framework and then introduce the trainable scheme via knowledge distillation.
4.1. IFRAgent pipline
The IFRAgent pipeline can be divided into 2 phases: the intention flow extraction phase and the deployment phase.
4.1.1. Intention Flow Extraction Phase
In the intention flow extraction phase, IFRAgent collects human demonstrations and analyzes these human demonstrations to extract the implicit intention flow and explicit intention flow of humans. For each user $u_{i}\in U=\{u_{1},u_{2},\dots,u_{n}\}$, we first collect human demonstrations comprising a set of queries $Q_{i}=\{q_{1},q_{2},\dots,q_{k}\}$ and initialize an empty user-level habit repository $h_{i}$. Each query $q_{j}\in Q_{i}$ is accompanied by a sequence of operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\dots,s_{p}\}$ provided by the user.
Given the tuple $(q_{j},S(u_{i},q_{j}))$, we first process it through an explicit intention flow agent $A_{e}$ to extract a SOP:
(1) $p_{j}=A_{e}(q_{j},S(u_{i},q_{j})).$
Concurrently, the query $q_{j}$ is encoded into a latent representation $\mathbf{l}_{j}$ using an embedding model $\phi$, such that $\mathbf{l}_{j}=\phi(q_{j})$. The pair $(\mathbf{l}_{j},p_{j})$ is stored for $u_{i}$ to facilitate retrieval during deployment.
Simultaneously, the tuple is processed through an implicit intention flow agent $A_{i}$ that incrementally updates the habit repository:
(2) $h_{i}\leftarrow h_{i}\cup\{A_{i}(h_{i},q_{j},S(u_{i},q_{j}))\}.$
Where $A_{i}$ learns latent behavioral patterns from interaction sequences. This dual-processing framework iterates over all queries in $Q_{i}$ until all human demonstrations are consumed, resulting in a comprehensive habit repository $h_{i}$ and retrievable explicit SOPs $\{(\mathbf{l}_{j},p_{j})\}$ for each user.
Table 2. Performance comparison of mobile-use agents with IFRAagent enhancements, categorized by open-source and close-source models. IFRAgent demonstrates improvements across nearly all metrics.
English users Chinese users
[…]
Open-source based mobile-use agents
OS-Atlas-7B-Pro 42.30 75.39 36.65 42.22 76.92 35.67
                    + IFRAgent 48.466.16↑ 80.985.59↑ 43.466.81↑ 51.979.75↑ 81.644.72↑ 47.8212.15↑
UI-TARS-7B-SFT 43.11 69.26 37.28 44.86 75.00 36.86
                    + IFRAgent 44.481.37↑ 72.573.31↑ 40.693.41↑ 51.046.18↑ 75.860.86↑ 48.7011.84↑
UI-TARS-7B-DPO 41.12 71.11 34.96 45.07 70.74 37.19
                    + IFRAgent 41.580.46↑ 70.910.20↓ 37.732.77↑ 46.451.38↑ 73.732.99↑ 43.466.27↑
UI-TARS-1.5-7B 40.19 72.06 34.02 46.22 76.42 39.18
                    + IFRAgent 42.782.59↑ 72.720.66↑ 39.265.24↑ 49.803.58↑ 77.010.59↑ 46.757.57↑
Qwen2.5-VL-7B 12.29 16.13 11.80 19.29 23.52 18.13
                    + IFRAgent 30.5718.28↑ 40.6724.54↑ 27.7015.90↑ 38.2718.98↑ 47.2723.75↑ 36.1318.00↑
GUI-owl-7B 42.99 82.71 38.13 39.32 70.21 35.07
                    + IFRAgent 50.23↑7.24 84.20↑1.49 44.63↑6.50 44.22↑4.90 73.26↑3.05 38.49↑3.42
Qwen3-VL-8B 40.75 77.34 36.64 38.08 65.57 32.09
                    + IFRAgent 48.79↑8.04 83.88↑6.54 44.07↑7.43 43.01↑4.93 70.21↑4.64 37.36↑5.27
Close-source based mobile-use agents
GPT-4o + OCR model 35.63 74.75 32.05 37.13 77.42 31.18
                    + IFRAgent 40.554.92↑ 74.500.25↓ 36.023.97↑ 44.197.06↑ 78.060.64↑ 41.4010.22↑
GLM-4v + OCR model 2.54 57.94 1.76 3.11 72.97 2.47
                    + IFRAgent 3.210.67↑ 73.1415.20↑ 2.470.71↑ 4.050.94↑ 73.710.74↑ 3.030.56↑
Qwen-VL-max + OCR model 19.86 79.91 16.69 22.00 81.55 19.04
                    + IFRAgent 20.370.51↑ 81.821.91↑ 17.560.87↑ 24.042.04↑ 84.422.87↑ 21.342.30↑
4.1.2. Deployment Phase
In the deployment phase, when processing a user query $q$ from user $u_{i}$, we first encode it into vector $\mathbf{l}$ using the same embedding model $\phi$ as employed in the intention flow extraction phase (where $\mathbf{l}$ = $\phi(q)$). Then we match $\mathbf{l}$ against each explicit intention flow $(\mathbf{l}_{j},p_{j})$ of the user $u_{i}$ for RAG.
When the similarity exceeds a threshold $\tau$, we obtain the most similar query $q^{\prime}$ and its corresponding SOP $p^{\prime}$:
(3) $(q^{\prime},p^{\prime})=\begin{cases}(q_{j},p_{j})&\text{if }\exists j=\arg\max_{k}\text{sim}(\mathbf{l},\mathbf{l}_{k})\\ &\quad\land\ \text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau,\\ \varnothing&\text{otherwise.}\end{cases}$
The query $q^{\prime}$, its SOP $p^{\prime}$, and the query $q$ are used together as the prompt for few-shot learning and then fed into the SOP Extractor $\mathcal{E}$ to obtain the SOP $p$ corresponding to $q$:
(4) $p=\mathcal{E}(q,(q^{\prime},p^{\prime})).$
Next, the query $q$ and its corresponding SOP $p$ are combined with the user habit repository $h_{i}$ of user $u_{i}$ as input to the query writer $\mathcal{W}$ to generate a rewritten personalized query $\hat{q}$ and SOP $\hat{p}$ that align user-specific intention:
(5) $(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i}).$
Finally, the rewritten query $\hat{q}$, the rewritten SOP $\hat{p}$, and the current screenshot $s$ are provided as input to the mobile-use agent $\mathcal{F}$ to obtain the action $a$:
(6) $a=\mathcal{F}(\hat{q},\hat{p},s).$
4.2. Training via Knowledge Distillation
The SOP Extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are two key components of the IFRAgent during the deployment phase. Both $\mathcal{E}$ and $\mathcal{W}$ are models deployable on the edge side, thus lacking general knowledge about mobile operations. Since $\mathcal{E}$ has already compensated for the lack of general knowledge through few-shot learning with RAG, only $\mathcal{W}$ requires supervised fine-tuning to unleash its potential in query rewriting.
To warm up the trainable scheme query writer $\mathcal{W}$, we employ a data distillation method, allowing $\mathcal{W}$ to learn the query and SOP rewriting logic from large-scale models with strong general knowledge (e.g., GPT-4o). For the support dataset of MobileIAR, we used GPT-4o to replace the query writer $\mathcal{W}$ during the deployment phase, obtaining personalized rewritten queries $\hat{q}$ and personalized rewritten SOPs $\hat{p}$ for each query. Then, using these $\hat{q}$ and $\hat{p}$ as labels, we performed supervised fine-tuning on Qwen3-4B to warm up the query writer $\mathcal{W}$, enabling it to learn how to rewrite queries and SOPs using personalized user information. The training objective can be formulated as:
[…]
After the warm-up training, $\mathcal{W}$ acquires the capability to rewrite queries and SOPs in a user-specific manner.
5. Experiments
5.1. Experiments Setup
5.1.1. Baseline
Mobile-use agents can be categorized by their construction methods into open-source and closed-source agents. We selected five state-of-the-art open-source-based mobile-use agents and three closed-source-based mobile-use agents as baselines to experimentally validate the extent to which IFRAgent enhances mobile-use agents. The open-source based mobile-use agents include OS-Atlas-7B-Pro (Wu et al., 2025), UI-TARS-7B-SFT, UI-TARS-7B-DPO (Qin et al., 2025), UI-TARS-1.5-7B, Qwen2.5-VL-7B (Bai et al., 2025b), GUI-owl-7B (Ye et al., 2025) and Qwen3-VL-8B (Bai et al., 2025a) while the close-source based mobile-use agents include GPT-4o (OpenAI, 2023), GLM-4v (GLM et al., 2024), and Qwen-VL-max (Bai et al., 2023). Since GPT-4o, GLM-4v, and Qwen-VL-max inherently lack the capability to predict coordinates, we incorporated an OCR model composed of ResNet18 (He et al., 2016) and ConvNeXt-Tiny (Liu et al., 2022) to assist them in localization. Our main experiments were conducted on our collected dataset, MobileIAR. MobileIAR is the first benchmark for user-specific intent alignment testing in the field of mobile-use agents. In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP extractor and query rewriter are based on Qwen3-4B. And the embedding model is jina-embeddings-v2.
5.1.2. Metric
We consider two types of metrics: one measures the task completion capability of mobile-use agents, and the other measures the alignment level between mobile-use agents and human intentions. Following existing work on mobile-use agents (Zhang and Zhang, 2024; Ma et al., 2024; Wu et al., 2025; Qin et al., 2025; Cheng et al., 2025a), we report the step-wise success rate (SR) and action type accuracy (Type) to assess task completion. To measure the alignment level between mobile-use agents and human intentions, we report the intention alignment rate (IAR). In the MobileIAR dataset, we provide both human-intent-aligned actions and ground-truth action lists. The human-intent-aligned action is the single most aligned action with the user’s intent at each step, while ground-truth action lists are a set of possible actions that could help fulfill the user’s query in the current frame. For metric calculations, SR and Type consider the mobile-use agent’s action correct if it matches any of the ground-truth actions at the current step. In contrast, IAR requires the agent’s action to exactly match the human-intent-aligned action to be counted as correct.
5.2. Main Results
As shown in Table 2, we conducted extensive experiments on mobile-use agents constructed using eight different methods. Based on these results, we have the following key findings:
(i) IFRAgent can improve almost all mobile-use agents, with absolute improvements of SR by 5.50% and IAR by 6.56%, and relative improvements of SR by 24.20% and IAR by 28.80%. This indicates that IFRAgent can enhance both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. Moreover, these improvements are effective for both English and Chinese users.
(ii) Among general-domain models, Qwen2.5-VL-7B, Qwen3VL-8B and GPT-4o demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS. This is because models with broader general knowledge tend to have a more accurate and comprehensive understanding of human intentions. Specialized mobile-use agent base models, due to the extensive GUI operation knowledge learned during the post-training process, tend to forget some general world knowledge that is helpful for human intention recognition.
(iii) Building personalized mobile-use agents requires the model to possess basic instruction-following capabilities. We can observe that the Type metric of mobile-use agents built using GLM-4v and Qwen-VL-max is significantly lower than the SR metric, indicating that in most cases, they can determine that the current action type required is CLICK, but fail to utilize the localization information provided by the OCR model. Therefore, even with the addition of the IFRAgent module, the improvement for mobile-use agents built with GLM-4v and Qwen-VL-max is not significant, as IFRAgent does not provide localization information for the mobile-use agent.
(iv) Overall, for Chinese user scenarios, mobile-use agents perform better. This is because Chinese app designs are more standardized; furthermore, model series such as OS-Atlas-7B-pro and UI-TARS series have been more extensively trained on data from Chinese environments.
Overall, IFRAgent effectively enhances mobile-use agents by improving task completion and human intention understanding, particularly for general-domain models. Based on the above experimental results and findings, what determines the lower bound of the personalized capability of a mobile-use agent is general language understanding ability, while what determines the upper bound is its execution capability within the mobile-use domain itself.
6. Further Analysis
In this section, we assess the contribution of individual components, evaluate generalizability across datasets, analyze scaling behavior, and measure computational efficiency.
6.1. Ablation Study
Figure 5. Experimental results of ablation study.
The SOP extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are the two most critical components in IFRAgent. We conducted an ablation study on Qwen2.5-VL-7B, the model with the most significant improvement from IFRAgent. The experimental results are shown in Figure 5. It can be observed that the SOP extractor $\mathcal{E}$ alone cannot fully unleash the potential of human demonstrations; after personalized rewriting, all metrics still show improvement. On the other hand, the query writer $\mathcal{W}$ alone, which only rewrites abstract personalized queries without SOP teaching the agent how to act, can only bring slight improvements over the baseline. Therefore, the design of IFRAgent maximizes the potential mined from human demonstrations and is reasonable.
Figure 6. Experimental results of IFRAgent on model scale for Qwen2.5-VL-3B, Qwen2.5-VL-7B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. The improvement is more significant for medium-scale mobile-use agents.
6.2. Generalizability Analysis on Other Datasets
Table 3. Generalizability experiment on our modified user-specific OS-Kairos dataset.
[…]
OS-Atlas-7B-Pro 58.85 79.52 53.78
                    + IFRAgent 69.74 85.31 68.42
UI-TARS-1.5-7B 61.49 75.86 53.24
                    + IFRAgent 60.75 77.04 58.80
Qwen2.5-VL-7B 24.27 27.77 23.64
                    + IFRAgent 51.64 59.62 50.31
GPT-4o+OCR model 40.65 68.92 37.67
                    + IFRAgent 56.33 77.36 53.75
To validate the generalizability of the IFRAgent method, we adapt a portion of the OS-Kairos dataset (Cheng et al., 2025a). We simulate the extraction of trajectories across shop, video, and search scenarios in the training set of OS-Kairos as human demonstrations. Then, in the corresponding scenarios of the OS-Kairos test set, we sample user trajectories and manually annotate the user-intent-aligned actions and ground-truth action lists. For each frame of OS-Kairos data, we had different annotators label their own user-intent-aligned actions based on the screenshot and query. Simultaneously, following the same construction method as the MobileIAR dataset, we used a joint annotation approach with groups of three people to label the ground-truth action list based on the trajectories that had been annotated with user-intent-aligned actions.
The experimental results are presented in Table 3. IFRAgent continues to demonstrate improvements in both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios on our modified user-specific OS-Kairos dataset, proving the generalizability of IFRAgent.
At the same time, consistent with the conclusions in the main results, general-domain models such as Qwen2.5-VL-7B and GPT-4o still show better performance than specialized mobile-use agent base models like UI-TARS.
6.3. Comparison With Other Methods
Table 4. Comparison with SOP demonstration. IFRAgent can better abstract the user’s intention flow, thereby significantly enhancing the agent.
[…]
OS-Atlas-7B-Pro 42.26 76.24 36.11
           + LearnAct 49.26 77.84 44.61
           + SOP demonstration 50.08 79.65 43.72
           + IFRAgent 50.40 81.35 45.87
UI-TARS-1.5-7B 43.53 74.48 36.88
           + LearnAct 42.26 69.43 38.86
           + SOP demonstration 41.32 70.86 37.17
           + IFRAgent 46.67 75.10 43.41
Qwen2.5-VL-7B 16.17 20.22 15.31
           + LearnAct 21.88 29.16 20.15
           + SOP demonstration 22.64 28.68 21.11
           + IFRAgent 34.83 44.32 32.37
GPT-4o+OCR model 36.46 76.23 31.57
           + LearnAct 42.29 75.01 38.46
           + SOP demonstration 43.37 76.13 39.29
           + IFRAgent 42.57 76.47 39.00
6.3.1. Implementation Details
To validate the effectiveness of IFRAgent in recognizing intention flow from human demonstrations, we compared it with other methods that extract information from human demonstrations. We use a demoparser to extract the SOP from a set of user-specific human demonstrations. For each query, we matched it with the corresponding 1-shot SOP demonstration to enhance the prompt. We then selected four representative baselines on MobileIAR to compare this SOP demonstration method with IFRAgent. Finally, we rewrite all the queries into ambiguous instructions to facilitate testing for IAR. Additionally, we also compared our results with the experimental results of Learnact (Liu et al., 2025a).
6.3.2. Result Analysis
As shown in Table 4, the experimental results demonstrate that all demonstration learning methods can improve various metrics across different baselines. However, for almost all baselines, IFRAgent achieves more significant improvements compared to other methods.
For the mobile-use agent built with GPT-4o, the performance of IFRAgent and other methods is relatively close. This is because GPT-4o, as a foundation model with extensive world knowledge, can leverage its own capabilities to abstract the user’s implicit intent flow from the SOP demonstration to assist in task completion.
For most 7B-scale agents, whether general-domain models or specialized mobile-use agent base models, IFRAgent shows significantly better performance than other methods. This is because 7B-scale agents struggle to accurately abstract implicit intent flow from SOP demonstrations. In contrast, IFRAgent already incorporates the user’s explicit and implicit intent flows into the personalized rewritten query and personalized rewritten SOP, eliminating the need for the agent to perform further abstraction.
6.4. Scale Analysis
6.4.1. Model scale Analysis
To analyze the impact of IFRAgent on mobile-use agents with different parameter scales and simultaneously verify its generalizability across varying model sizes, we conducted experiments on Qwen2.5-VL-3B, Qwen2.5-VL-7B, and Qwen2.5-VL-72B, respectively. As shown in Figures 6, IFRAgent improves both task completion capability and alignment with human intentions for mobile-use agents of different model scales. The most significant improvements are observed in the 7B and 32B model scales. This is because the 3B-parameter mobile-use agent lacks sufficient instruction-following ability—even when provided with personalized queries and SOPs, the mobile-use agent still fails to comply. On the other hand, the 72B-parameter mobile-use agent already possesses strong general mobile operation capabilities, so the SR improvement is less pronounced. However, there remains a noticeable enhancement in the IAR for the 72B-parameter mobile-use agent. IFRAgent stimulates both the general task completion capability and the ability to align with human intentions in mobile-use agents with moderate parameter scales.
Figure 7. Experimental results of IFRAgent-enhanced model under varying numbers of demonstrations.
6.4.2. Demonstration Count Analysis
To analyze the impact of the number of demonstration queries and SOPs on the Extractor $\mathcal{E}$, we conducted a demonstration count analysis on Qwen2.5-VL-7B. We input varying numbers of $(q^{\prime},p^{\prime})$ pairs into $\mathcal{E}$ for SOP extraction. The experimental results are shown in Figure 7. The results show that while 1-shot brings a significant improvement over 0-shot, increasing to 2-shot and 3-shot does not lead to further substantial gains across metrics and may even introduce irrelevant information that interferes with $\mathcal{E}$, causing performance degradation. On one hand, this indicates that the intent flow extracted by IFRAgent from human demonstrations can indeed help mobile-use agents align with human intent. On the other hand, increasing the number of demonstrations entails higher computational overhead. Therefore, using 1-shot for SOP extraction in IFRAgent is a reasonable choice.
6.5. Time Cost of IFRAgent
To evaluate the time cost of IFRAgent, we conducted experiments by adding IFRAgent on top of the base models. As shown in Figure 8, we performed experiments on both open-source models (7B and 72B) and larger-scale closed-source models. Since both query rewriting and SOP generation are handled by Qwen3-4B and each query is processed only once, the additional time overhead is minimal. This indicates that IFRAgent not only improves the alignment between mobile-use agents and human intent, but is also highly efficient.
Figure 8. Comparison of inference time with and without IFRAgent across different model scales.
7. Conclusion
This study focuses on two major challenges in building personalized mobile-use agents: the lack of alignment evaluation benchmarks and the impracticality of per-user fine-tuning. For these challenges, this study collects and open-source the MobileIAR dataset, a user-specific dataset designed to assess the intention alignment rate between mobile-use agents and humans. This study then proposes IFRAgent, a plug-and-play framework based on intention flow recognition from human demonstrations, which enhances the alignment between agents and human intentions by leveraging human demonstrations to rewrite user instructions into personalized queries and SOPs. The extensive experiment results show that IFRAgent improves both the task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. This study also provides various analyses to offer valuable insights for building personalized mobile-use agents.